OpenStudy (anonymous):

What are the intercepts of the equation? Graph the equation. 5x+3y=15

jimthompson5910 (jim_thompson5910):

To get the y-intercept, you plug in x = 0 5x+3y=15 5*0+3y=15 0+3y=15 3y=15 y = 15/3 y = 5 This means the y-intercept is (0,5)

jimthompson5910 (jim_thompson5910):

To get the x-intercept, plug in y = 0 and solve for x. Tell me what you get.

OpenStudy (anonymous):

x intercept would be (5,0)?

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

5x+3y=15 5x+3*0=15 x = ???

OpenStudy (anonymous):

that would make it (3,0)

jimthompson5910 (jim_thompson5910):

so you now have the two points (3,0) and (0,5)

jimthompson5910 (jim_thompson5910):

plot them both on the same coordinate axis then draw a straight line through them

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

you're welcome